Suppose your utility function is u(x)-VX Consider the following two games: 1. You have a 1/4 chance of winning 25 and a 3/4 chance of winning 1 2. You have a 2/3 chance of winning 7 and a 1/3 chance of winning 4 What is the expected value and the expected utility of these games?
Answer #1

1) u(x) =\sqrt{}x

expected value=\sum x* P(x)

=1/4 *25 + 3/4*1 = 7

expected utility =\sumP(x) * u(x)

=1/4 * \sqrt{} 25 + 3/4 * \sqrt{} 1 = 2

___________________________________________________

2)

expected value=\sum x* P(x)

=2/3 * 7 + 1/3 * 4

=18/3

expected utility =\sumP(x) * u(x)

2/3 * \sqrt{} 7 + 1/3 *\sqrt{}4

=(2/3) (\sqrt{}7 + 1)

=1.2244